# Hyperparameter Importance Analysis for Multi-Objective AutoML

### v2

@@ Abstract
Hyperparameter optimization plays a pivotal role in enhancing the predictive performance and generalization capabilities of ML models. However, in many applications, we do not only care about predictive performance but also about objectives such as inference time, memory, or energy consumption. In such MOO scenarios, determining the importance of hyperparameters poses a significant challenge due to the complex interplay between the conflicting objectives. In this paper, we propose the first method for assessing the importance of hyperparameters in the context of multi-objective hyperparameter optimization. Our approach leverages surrogate-based hyperparameter importance (HPI) measures, i.e. fANOVA and ablation paths, to provide insights into the impact of hyperparameters on the optimization objectives. Specifically, we compute the a-priori scalarization of the objectives and determine the importance of the hyperparameters for different objective tradeoffs. Through extensive empirical evaluations on diverse benchmark datasets with three different objectives paired with accuracy, namely time, demographic parity, and energy consumption, we demonstrate the effectiveness and robustness of our proposed method. Our findings not only offer valuable guidance for hyperparameter tuning in MOO tasks but also contribute to advancing the understanding of HPI in complex optimization scenarios.
@@ 1 Introduction
The selection of appropriate hyperparameter configurations significantly impacts a model’s ability to capture underlying patterns in the data and produce accurate predictions. Optimizing hyperparameters is one of the main focus areas of Automated Machine Learning (AutoML) [21, 7]. Given a use case and a model, it is usually not known which hyperparameters are worth tuning to achieve a good performance. Providing insights into this (e.g., [36]) is valuable since it allows the design of better configuration spaces and gives a better understanding of the learning dynamics of machine learnirng (ML) algorithms. Hyperparameter Importance (HPI) offers a systematic method to gain insights into the influence of hyperparameters on the model’s performance. Understanding and optimizing hyperparameters are crucial steps in building effective ML models. They allow us to fine-tune our algorithms, improve performance, and achieve better generalization.
So far, most efforts in hyperparameter optimization (HPO) have been focused on single-objective optimization, mainly targeting the predictive performance of models. Recently, there has been a trend towards multi-objective HPO [17, 3, 13, 15, 6, 22], which allows the consideration of several objectives, including fairness, memory consumption, training time, inference time, and energy consumption. In a multi-objective optimization (MOO) scenario, where we optimize conflicting objectives simultaneously, it becomes not only more challenging to approximate the unknown Pareto front of non-dominated configurations but also to discern the relative significance of individual hyperparameters.
@@
Conventional methods for assessing HPI rely on univariate sensitivity analysis techniques, like variance-based methods [20] or partial dependence plots [28]. For instance, techniques like fANOVA have been widely used to decompose the variance of model performance into contributions from individual hyperparameters [20]. While effective for single-objective optimization, these methods may not capture the intricate interactions among multiple objectives in multi-objective settings. So far, there are no methods that assess the importance of hyperparameters within the context of MOO.
@@
In this paper, we propose a novel method called weighted multi-objective hyperparameter importance (MO-HPI). Although our approach is in principle usable with any surrogate-based HPI measure, we specifically focus on two widely-used methods, fANOVA [20] and ablation path analysis [14, 4]. Our methodology involves training surrogate models, e.g., random forests, with the hyperparameter configuration data and the respective objective results. We use weighted sums of the objectives as target variables, allowing us to effectively capture the change in importance from one objective to another. Figure 1 shows the concept of MO-fANOVA and Figure 2 gives an overview of the idea for the MO-ablation path analysis.
This work aims to propose a method to analyze MOO results and demonstrate that it is reasonable in the context of ML. To reach that goal, we pose the following research question: How to convert surrogate-based HPI methods to meaningfully measure the importance of hyperparameters within multi-objective optimization? We further frame the following subquestions: Are the results intuitively correct? How do the proposed methods compare? Can we gain new insights into the impact of hyperparameters on the optimization process that previous methods could not provide?
To answer our research questions and validate the efficacy of our proposed method, we conduct three empirical evaluations on the well-known benchmark datasets MNIST [9], Adult Census Income [2] and CIFAR10 [24]. To show that the approach can be applied to diverse objectives, we consider three objectives paired with accuracy: training time, demographic parity, and energy consumption, and discuss the results for each of them accordingly.
@@ 2 Related Work
Another method is ablation path analysis [14, 4], which compares the default configuration and the optimized configuration to measure hyperparameter contributions. It creates an ablation path from the default configuration to a target configuration by changing the hyperparameter with the largest increase in performance in each iteration. We note that these ablation paths do not follow traditional ablation studies in which only a single hyperparameter is changed and all others are fixed; but this allows a full path through the configuration space from the default configuration to a target configuration. While there is only one path in a single-objective scenario, there are potentially many in the MOO setting (cf. Figure 2).
fANOVA [20] identifies the importance of individual hyperparameters and interactions among them. For each hyperparameter, it measures how it contributes to the variance in performance. This is done, e.g., by training a random forest as a surrogate model and subsequently decomposing the variance of each tree into contributions to each subset of hyperparameters. Based on these methods, several works have been published. For example, van Rijn and Hutter [36] applied HPI across datasets. PED-ANOVA [38] improved fANOVA using Pearson divergence to work better on arbitrary subspaces of the search space, e.g. the subspace of the top-performing configurations. This paper contributes to this body of work by extending the methodology to MOO.
@@ 3 Weighted HPI for Multiple Objectives
Our main idea is simple and intriguing given the insight that we are interested in the HPI for different objectives tradeoffs. The main challenge lies in (i) how to map this to different HPI methods, (ii) how to compute this efficiently, (iii) how to visualize this, and (iv) how to interpret this. Challenges (i)-(iii) are described in this section and (iv) follows as part of the following experiment and discussion sections. Overall, our work employs a framework for assessing the influence of hyperparameters on multiple objectives using any surrogate-based HPI analysis methods, allowing it to be efficient in computing analyses. Given a performance dataset obtained by an MO-HPO optimizer, surrogate models are trained to calculate the HPI. In the following, the steps will be explained in more detail. W.l.o.g. we discuss our approaches for two competing objectives; the approach can be straightforwardly extended to more objectives.
@@
The performance dataset should contain model configurations with their performance on several objectives. First, each of the objectives is normalized using the min-max normalization. Next, the configuration data is converted to numerical values to deal with categorical data and nan values.
@@ 3.2 Weighting Scheme
The objectives are weighted by the tradeoffs of the optimized configurations. Those are the configurations where the objective values are on the Pareto front. The configurations on the Pareto front are the set of non-dominated solutions. To retrieve the corresponding weighting, only Pareto-efficient objective pairs $(o_{1},o_{2})$ of the non-dominated solutions are considered. Subsequently, the normalized objective pairs are scaled to sum up to 1 by adding the pair up and dividing each value by the total, to be used as weighting.
@@ 3.3 Multi-Objective fANOVA
1:$(\vec{O_{1}},\vec{O_{2}})\leftarrow$ normalized objective values
2:$X\leftarrow$ encoded configuration data
3:$\vec{W}\leftarrow$ is_pareto_efficient($(\vec{O_{1}},\vec{O_{2}})$)
4:for each $\vec{w}$ in $\vec{W}$ do
5:     $Y_{w}\leftarrow w_{1}\cdot\vec{O_{1}}+w_{2}\cdot\vec{O_{2}}$
6:     $\mathcal{S}\leftarrow$ surrogate model trained with $(X,Y_{w})$
7:     Calculate fANOVA importances with $\mathcal{S}$ following [20]
8:end for
A single target variable (Y) is created for each weighting $(w_{1},w_{2})$ by summing the normalized weighted objectives values $(\vec{O_{1}},\vec{O_{2}})$:
$Y_{w}=w_{1}\cdot\vec{O_{1}}+w_{2}\cdot\vec{O_{2}}$ (1)
@@ 3.4 Multi-Objective Ablation Path Analysis
Input: evaluated configurations $\lambda\in\vec{\Lambda}$, corresponding evaluated objective values $(\vec{O_{1}},\vec{O_{2}})$, hyperparameters $hps$
1:$(\vec{O_{1}},\vec{O_{2}})\leftarrow$ objective values
2:$\vec{\Lambda}\leftarrow$ encoded configuration data
3:$\vec{W}\leftarrow$ is_pareto_efficient($(\vec{O_{1}},\vec{O_{2}})$)
4:for each $obj$ in $(\vec{O_{1}},\vec{O_{2}})$ do $\triangleright$ Not normalized
5:     $\vec{\mathcal{S}}[{obj}]\leftarrow$ random forest trained with $\vec{\Lambda},Y_{\text{obj}}$
6:end for
7:for each $\vec{w}$ in $\vec{W}$ do
8:     $\lambda_{\text{best}}\in\underset{\vec{\Lambda}}{\operatorname*{arg\,min}}(w_{1}\cdot\vec{O_{1}}+w_{2}\cdot\vec{O_{2}})$
9:     $r_{\text{previous}}\leftarrow\sum(\vec{w}\times\vec{\mathcal{S}}(\lambda_{\text{default}}))$
10:     $r_{\text{min}}\leftarrow r_{\text{previous}}$
@@
13:              $\lambda_{\text{default}}[hp]\leftarrow\lambda_{\text{best}}[hp]$
14:              for each $\mathcal{S}_{\text{obj}}$ in $\vec{\mathcal{S}}$ do
15:                  $\vec{r}[obj]\leftarrow$ normalized $\mathcal{S}_{\text{obj}}(\lambda_{\text{default}})$
16:              end for
17:              $r_{\text{total}}\leftarrow\sum(\vec{w}\times\vec{r})$triangleright Scalarized performance
18:              if $r_{\text{total}}<r_{\text{min}}$ then
19:                  $hp_{\text{min}}\leftarrow hp$triangleright Most important hyperparameter
20:                  $r_{\text{min}}\leftarrow r_{\text{total}}$
21:              end if
22:              reset $\lambda_{\text{default}}$
23:         end for
24:         Add $hp_{\text{min}}$ with $r_{\text{min}}$ to ablation path
25:         Remove $hp_{\text{min}}$ from $hps$
26:     end while
27:end for
We propose to extend the ablation path analysis for multiple objectives as follows, see also Algorithm 2. For both objectives, a surrogate model is trained with the encoded configuration data and the objectives before normalization (line 4). For each weighting, the ablation path is calculated (starting from line 7). The incumbent $\lambda_{best}$ is the corresponding configuration to the minimum of the scalarized target variable $Y_{w}$ of Equation 1 (line 8). Starting from the default configuration, each hyperparameter value is changed individually to the value of the incumbent configuration, and the performance is estimated with the trained surrogate models (lines 12-16). The results of the models are normalized with the same normalization as the respective objective values, weighted, and summed up as performance (line 15 and 17). The hyperparameter value leading to the highest difference in performance (toward the optimization objective) will be changed in the configuration (lines 18-21 and line 24). The ablation is repeated with the altered configuration and the remaining hyperparameters until the incumbent configuration is reached (line 11). The difference in performance is recorded for each weighting.
@@ 4 Experiments
Using our MO-HPI approach, we evaluated three HPO problems with different data, models, and objectives. The overview can be seen in Table 1. We chose rather simple benchmarks and models so that the results of the MO-HPI approach are easier to validate by common knowledge.
@@
fairness Adult Census MLP 1-Accuracy Demographic Parity
@@ 4.1 General Setup
The basic setup was the same for all experiments. We used SMAC [18, 26] as one of the state-of-the-art HPO tools [12] to perform HPO. The multi-objective algorithm ParEGO [23] with Random Online Aggressive Racing (ROAR) [18] and 1000 trials was used for optimization. ROAR randomly selects a configuration from the hyperparameter space and only keeps track of the top 10 non-dominated configurations. We note that our primary goal was not to achieve the best possible multi-objective performance but to use a sufficient amount of performance data generated by an HPO tool. Since Moosbauer et al [28] showed post-hoc analysis of HPO runs can be biased by too strong exploitation, we focus here on unbiased randomly generated data. For both HPI methods, we trained random forests with 100 trees as surrogate models.

### v3

@@ Abstract
Hyperparameter optimization plays a pivotal role in enhancing the predictive performance and generalization capabilities of ML models. However, in many applications, we do not only care about predictive performance but also about additional objectives such as inference time, memory, or energy consumption. In such multi-objective scenarios, determining the importance of hyperparameters poses a significant challenge due to the complex interplay between the conflicting objectives. In this paper, we propose the first method for assessing the importance of hyperparameters in multi-objective hyperparameter optimization. Our approach leverages surrogate-based hyperparameter importance measures, i.e., fANOVA and ablation paths, to provide insights into the impact of hyperparameters on the optimization objectives. Specifically, we compute the a-priori scalarization of the objectives and determine the importance of the hyperparameters for different objective tradeoffs. Through extensive empirical evaluations on diverse benchmark datasets with three different objective pairs, each combined with accuracy, namely time, demographic parity loss, and energy consumption, we demonstrate the effectiveness and robustness of our proposed method. Our findings not only offer valuable guidance for hyperparameter tuning in multi-objective optimization tasks but also contribute to advancing the understanding of hyperparameter importance in complex optimization scenarios.
@@ 1 Introduction
The selection of appropriate hyperparameter configurations significantly impacts a model’s ability to capture underlying patterns in the data and produce accurate predictions. Optimizing hyperparameters is one of the main focus areas of Automated Machine Learning (AutoML) [22, 7]. Given a use case and a model, it is usually unknown which hyperparameters are worth tuning to achieve a good performance. Providing insights into this (e.g., [34, 42, 31]) is valuable since it allows the design of better configuration spaces and gives a better understanding of the learning dynamics of Machine Learning (ML) algorithms. HyperParameter Importance (HPI) offers a systematic method to gain insights into the influence of hyperparameters on the model’s performance. Understanding and optimizing hyperparameters are crucial steps in building effective ML models. They allow us to fine-tune our algorithms, improve performance, and achieve better generalization.
@@
So far, most efforts in hyperparameter optimization (HPO) have been focused on single-objective optimization, mainly targeting the predictive performance of models. Recently, there has been a trend towards multi-objective HPO [18, 3, 13, 15, 6, 23], which allows the consideration of several objectives, including fairness, memory consumption, training time, inference time, and energy consumption. In a Multi-Objective Optimization (MOO) scenario, where we optimize conflicting objectives simultaneously, it becomes more challenging to approximate the unknown Pareto front and to discern the relative significance of individual hyperparameters. The configurations on the Pareto front are the set of non-dominated solutions.
@@
Conventional methods for assessing HPI in single-objective scenarios rely on univariate sensitivity analysis techniques, like variance-based methods [21] or partial dependence plots [30]. For instance, techniques like fANOVA have been widely used to decompose the variance of model performance into contributions from individual hyperparameters [21]. While effective for single-objective optimization, these methods may not capture the intricate interactions among multiple objectives in multi-objective settings. So far, no methods have assessed the importance of hyperparameters within the context of MOO.
In this paper, we propose a novel method called weighted multi-objective HyperParameter Importance (MO-HPI). Although our approach is, in principle, usable with any surrogate-based HPI measure, we specifically focus on two widely-used methods, fANOVA [21] and ablation path analysis [14, 4]. Our methodology involves training surrogate models, e.g., random forests, with the hyperparameter configuration data and the respective objective results. For better visualiziability and interpretability, we focus on bi-objective scenarios, but in principle, our approach can be used for any number of objectives. We use weighted sums of the objectives as target variables, allowing us to effectively capture the change in importance from one objective to another. Figure 1 shows the concept of MO-fANOVA, and Figure 2 the idea for the MO-ablation path analysis.
To analyze MO-HPI and demonstrate that it is reasonable in the context of ML, we pose the following research question: How to convert surrogate-based HPI methods to meaningfully measure the importance of hyperparameters within multi-objective optimization? We further frame the following subquestions: Are the results intuitively correct? How do the proposed methods compare? Can we gain new insights into the impact of hyperparameters on the optimization process that previous methods could not provide?
To answer our research questions and validate the efficacy of our proposed method, we conduct three empirical evaluations on the well-known benchmark datasets MNIST [9], Adult Census Income [2] and CIFAR10 [25]. To show that the approach can be applied to diverse objectives, we consider three pairs of objectives, each combined with accuracy: training time, demographic parity loss, and energy consumption, and discuss the results accordingly.
@@ 2 Related Work
Another method is ablation path analysis [14, 4], which compares the default and optimized configuration to measure hyperparameter contributions. It creates an ablation path from the default configuration to a target configuration by changing the hyperparameter with the largest increase in performance in each iteration. We note that these ablation paths do not follow traditional ablation studies in which only a single hyperparameter is changed, and all others are fixed; however, this allows a full path through the configuration space from the default configuration to a target configuration. While there is only one path in a single-objective scenario, potentially, there are many in the MOO setting (cf. Figure 2).
fANOVA [21] identifies the importance of individual hyperparameters and interactions among them. For each hyperparameter, it measures how it contributes to the variance in performance. This is done, e.g., by training a random forest as a surrogate model and subsequently decomposing the variance of each tree into contributions to each subset of hyperparameters. Based on these methods, several works have been published. For example, van Rijn and Hutter [42], Probst et al. [34] and Moussa et al. [31] considered HPI across datasets. PED-ANOVA [44] improved fANOVA using Pearson divergence to work better on arbitrary subspaces of the search space, e.g., the subspace of the top-performing configurations. This paper contributes to this body of work by extending the methodology to MOO.
@@ 3 Weighted HPI for Multiple Objectives
Our main idea is simple and intriguing, given the insight that we are interested in the HPI for different objective tradeoffs. The main challenge lies in (i) how to map this to different HPI methods, (ii) how to compute this efficiently, (iii) how to visualize this, and (iv) how to interpret this. Challenges (i)-(iii) are described in this section, and (iv) follows as part of the following experiment and discussion sections. Overall, our work employs a framework for assessing the influence of hyperparameters on multiple objectives using any surrogate-based HPI analysis methods, allowing it to be efficient in computing analyses. Given a performance meta-dataset obtained by a MO-HPO optimizer, we train surrogate models to predict the objectives on unevaluated configurations; this is then used to calculate the HPI. The steps will be explained in more detail in the following section. While we discuss our approaches for two competing objectives, the approach can be straightforwardly extended to more objectives. The implementation is available on GitHub at https://github.com/automl/hpi_for_mo_automl.
In the following, we use the following notation. $(O_{1},O_{2})$ denote two objective values. $(\tilde{O_{1}},\tilde{O_{2}})$ are the normalized versions of them. We use $\vec{W}$ to denote a vector of weights weighting the objectives. If we sum the weighted objectives, we obtain $Y_{w}$. All the evaluated configurations are collected in $\Lambda$, where $\lambda\in\Lambda$ denotes one evaluated configuration. $hps$ are hyperparameter names and $hp_{\text{min}}$ is the most important one. To get the predicted (weighted) objective values $r$ of any configuration, we train two surrogate models $\mathcal{S}_{obj}$, each trained with $\Lambda$ and $O_{\text{obj}}$ for the ablation path analysis, weighted after prediction. For fANOVA, we train one surrogate model per weighting $\mathcal{S}_{w}$ with $\Lambda$ and $Y_{w}$ (weighted before training).
@@
The performance meta-dataset contains model configurations with their performance on several objectives. That means a table where each row contains hyperparameter configurations with the respective performance of all measured objectives. First, each objective is normalized using the min-max normalization across all evaluated configurations. Normalization is done to scale objectives, such as time, so that all objectives live on the same scale and thus are comparable in their magnitude. Next, the configuration data is converted to numerical values to deal with categorical data and NaN values. NaN values could occur because of hierarchical structures based on conditional hyperparameters, i.e., hyperparameters that are only part of the configuration if a certain value of another hyperparameter is chosen.
@@ 3.2 Weighting Scheme
Each configuration on the Pareto front implicitly refers to one weighted tradeoff of the objectives. Only Pareto-efficient objective pairs $(o_{1},o_{2})$ are considered to retrieve the corresponding weighting. Subsequently, the normalized objective pairs are scaled to sum up to 1 by adding each pair up and dividing each of the two values by that sum to be used as weighting ($\tilde{o_{1}}=\frac{o_{1}}{o_{1}+o_{2}}$).
@@ 3.3 Multi-Objective fANOVA
Input: evaluated and encoded configurations $\lambda\in\Lambda$, corresponding evaluated objective values $(O_{1},O_{2})$
1:$(\tilde{O_{1}},\tilde{O_{2}})\leftarrow$ normalized objective values
2:$\vec{W}\leftarrow$ is_pareto_efficient($(\tilde{O_{1}},\tilde{O_{2}})$)
3:for each $\vec{w}$ in $\vec{W}$ do
4:     $Y_{w}\leftarrow w_{1}\cdot\tilde{O_{1}}+w_{2}\cdot\tilde{O_{2}}$
5:     $\mathcal{S}_{w}\leftarrow$ surrogate model trained with $(\Lambda,Y_{w})$
6:     Calculate fANOVA importances with $\mathcal{S}_{w}$, e.g. following [21]
7:end for
A single target variable (Y) is created for each weighting $(w_{1},w_{2})$ derived from the evaluated configurations $\lambda\in\Lambda$ on the Pareto front by summing the weighted objectives values $(O_{1},O_{2})$:
$Y_{w}=w_{1}\cdot O_{1}+w_{2}\cdot O_{2}$ (1)
@@ 3.4 Multi-Objective Ablation Path Analysis
Input: evaluated and encoded configurations $\lambda\in\Lambda$, corresponding evaluated objective values $(O_{1},O_{2})$, hyperparameters $hps$
1:$\vec{W}\leftarrow$ is_pareto_efficient$((O_{1},O_{2}))$
2:for each $obj$ in $(O_{1},O_{2})$ do $\triangleright$ Not normalized
3:     $\mathcal{S}_{obj}\leftarrow$ surrogate model trained with $\Lambda,O_{\text{obj}}$
4:end for
5:for each $\vec{w}$ in $\vec{W}$ do
6:     $\lambda_{\text{best}}\leftarrow\underset{\Lambda}{\operatorname*{arg\,min}}(w_{1}\cdot O_{1}+w_{2}\cdot O_{2})$
7:     $r_{\text{previous}}\leftarrow w_{1}\cdot\mathcal{\tilde{S}}_{1}(\lambda_{\text{default}})+w_{2}\cdot\mathcal{\tilde{S}}_{2}(\lambda_{\text{default}}))$
8:     $r_{\text{min}}\leftarrow r_{\text{previous}}$
9:     $\lambda_{\text{previous}}\leftarrow\lambda_{\text{default}}$
10:     $\lambda_{\text{current}}\leftarrow\lambda_{\text{previous}}$
@@
13:              $\lambda_{\text{current}}[hp]\leftarrow\lambda_{\text{best}}[hp]$
14:              $r\leftarrow w_{1}\cdot\mathcal{\tilde{S}}_{1}(\lambda_{\text{current}})+w_{2}\cdot\mathcal{\tilde{S}}_{2}(\lambda_{\text{current}}))$
15:              if $r<r_{\text{min}}$ then
16:                  $hp_{\text{min}}\leftarrow hp$triangleright Most important hyperparameter
17:                  $r_{\text{min}}\leftarrow r_{\text{total}}$
18:              end if
19:              $\lambda_{\text{current}}\leftarrow\lambda_{\text{previous}}$
20:         end for
21:         Add $hp_{\text{min}}$ with $r_{\text{min}}$ to ablation path
22:         $\lambda_{\text{previous}}[hp_{\text{min}}]=\lambda_{\text{best}}[hp_{\text{min}}]$
23:         Remove $hp_{\text{min}}$ from $hps$
24:     end while
25:end for
We propose to extend the ablation path analysis for multiple objectives as follows; see also Algorithm 2. For both objectives, a surrogate model is trained with the encoded configuration data and the objectives before normalization (Line 2). For each weighting, the ablation path is calculated (starting from line 5). The incumbent $\lambda_{\text{best}}$ is the corresponding configuration to the minimum of the scalarized target variable $Y_{w}$ of Equation 1 (Line 6). Starting from the default configuration, each hyperparameter value is changed individually to the value of the incumbent configuration, and the performance is estimated with the trained surrogate models (Lines 12-20). The results of the models are normalized with the same normalization as the respective objective values, weighted, and summed up as performance (Line 14). The hyperparameter value leading to the highest difference in performance (toward the optimization objective) will be changed in the configuration (Lines 15-18 and line 22). The ablation is repeated with the altered configuration $\lambda_{\text{current}}$ and the remaining hyperparameters until the incumbent configuration is reached (Line 11). The difference in performance is recorded for each weighting.
@@ 4 Experiments
Using our multi-objective HyperParameter Importance (MO-HPI) approach, we evaluated three HPO problems with different data, models, and objectives. The overview can be seen in Table 1. We chose rather simple benchmarks and models so that the results of the MO-HPI approach are easier to validate by common knowledge.
@@
fairness Adult Census MLP 1-Accuracy Demographic Parity Loss
@@
Figure 3: Results for the time experiment. The Pareto front is on the left (error vs. training time in seconds), with the red dot being the default performance. The MO-fANOVA results are in the middle, and the MO-ablation path analysis is on the right. The x-axis corresponds to the weighting of the minimum error objective.
@@ 4.1 General Setup
The basic setup was the same for all experiments. We used SMAC [19, 27] as one of the state-of-the-art HPO tools [12] to perform HPO. The multi-objective algorithm ParEGO [24] with Random Online Aggressive Racing (ROAR) [19] and 1000 trials was used for optimization. ROAR randomly selects a configuration from the hyperparameter space and only keeps track of the top 10 non-dominated configurations. We note that our primary goal was not to achieve the best possible multi-objective performance but to use a sufficient amount of performance data generated by an HPO tool. Since Moosbauer et al. [30] showed that post hoc analysis of HPO runs can be biased by too strong exploitation, we focus here on unbiased randomly generated data. For both HPI methods, we trained random forests with 100 trees as surrogate models. Note that reproducing our experiments will lead to slightly different results for the time and energy experiment, as those variables depend on factors such as the hardware.